OpenStudy (anonymous):

What is a possible first step to eliminate a variable in the following system? 2x – 3y = –1 3x – 4y = –3 a. Multiply the first equation by –3 and the second equation by 2. b.Multiply the first equation by 3 and the second equation by –2. c.Multiply the first equation by –4 and the second equation by 3. d.All of the above.

OpenStudy (turingtest):

you want to figure out which of these actions will get the coefficients of at least one of the variables to be equal and opposite for example ax+by=c -ax+py=q we can add these: __________ (b+p)y=c+q

OpenStudy (turingtest):

notice that eliminated the x variable because the first had coefficient a and the second had -a

OpenStudy (turingtest):

so try each out what do you get when you a. Multiply the first equation by –3 and the second equation by 2. ?

OpenStudy (anonymous):

i got a to

OpenStudy (turingtest):

a "to" ?

OpenStudy (anonymous):

im answer that i got was a

OpenStudy (turingtest):

ok, why do you think it's a ?

OpenStudy (turingtest):

I was just asking you to try it out ans see what you get try doing what (a) says, and write what you get after those operations what are the two equations now?

OpenStudy (turingtest):

I'm in no way suggesting that a is the right answer

OpenStudy (anonymous):

is it c

OpenStudy (turingtest):

It kinda seems like you're just guessing unless you can tell me why you think a particular answer is right, so I'm not going to confirm or deny anything yet. Please explain your reasoning and hopefully I can help you better

OpenStudy (turingtest):

what do you get by multiplying the first equation by -3 ?